2 A particle travels in a straight line a distance of 2 m in a time of 0.01 seconds. It is known that the particle accelerates from rest with constant acceleration. What is the acceleration of the particle?

2

A particle travels in a straight line a distance of 2 m in a time of 0.01 seconds. It is known that the particle accelerates from rest with constant acceleration. What is the acceleration of the particle?

Grade:12

1 Answers

Nirmal Singh.
10 years ago

we know that s = ut + 0.5 a t^2
u = 0 , s = 2 m, t =0.01 s
2 = 0*0.01 + 0.5*a*(0.01)^2
a = 2 / (0.5 x 10^-4)
a = 4 x x10^4 m /s ^2
Thanks & Regards,
Nirmal Singh
Askiitians Faculty